Tell that is equatrials triamgle if know apex that are A(-2,-2) B(2,6) and C(6,-2)

Grade:12

1 Answers

Aditya Gupta
4 years ago
hello enis, your ques is very unclear as it has several typos and errors. kindly repost with pic attached. i think you are asking whether the triangle with vertices A(-2,-2) B(2,6) and C(6,-2) is equilateral or not.
so obviously AB^2= 8^2+4^2
and BC^2= 4^2+8^2
and AC^2= 8^2
hence triangle is isosceles but not equilateral.
kindly approve :=)